molly is only girl in Wilson family and the eldest of her parent's 7 children. the children's ages run in 4-year steps, with molly being 4 times as old as Sammy, the "baby" of the family. How old is Sammy?
Answer by rajagopalan(174)

Molly is the eldest
her parent's 7 children.
the children's ages run in 4-year steps,
So minimum age for Molly =7x4=28
Then 7 Children's ages are..4,8,12,16,20,24 and 28.
This gives age ratio 28/4 =7 which is not acceptable.
so try 8 as baby age
we get 8,12,16,20,24,28 and 31.
Age ratio = 32/8 =4
Ok
Answer : Sammys age is 8.